OpenStudy (jsilvers02):

lim x>1 (lnx/sinpix)

OpenStudy (jsilvers02):

Why is it negative pi? I understand the rest dealing with L'H rule. However, I am not finding the -pi part and just getting (1/x)/piCospiX which becomes (1/1) /pi*-1*1 which should be when simplified 1/-pi. So how from here is the answer simply 1/pi

OpenStudy (anonymous):

One method is by series. Realizing lim x>1 (lnx/sinpix) is the same thing as lim x>0 [ ln(1-x) / sin(pi(1-x)) ], and that sin(pi(1-x)) = sin(pi(x)) we get the new limit: lim x>0 [ ln(1-x) / sin(pi*x) ] we may then use taylor expansions of both functions: ln(1-x) = x + (1/2)x^2 + (1/3)x^3 + ... sin(pi*x) = pi*x - (1/3!)(pi*x)^3 + (1/5!)(pi*x)^5 - ... when placed in fractional form we may divide out an x, leaving 1 and a series on top, pi and some series on the bottom, as x -> 0 the series left on top and bottom obviously approach 0, leaving 1 and pi, the solution being 1/pi.

OpenStudy (jsilvers02):

Is there another method, aside from series?